Solve the following linear equations using the Elimination method and The Substitution method
1) X + 3 Y = 2
2X + 5 Y = 3

2) 3X + 4 Y = 12
6X + 2 Y = 6



Answer by ikleyn(52786)

.
Solve the following linear equations using the Elimination method and The Substitution method
1) X + 3 Y = 2
2X + 5 Y = 3
2) 3X + 4 Y = 12
6X + 2 Y = 6
~~~~~~~~~~~~~~~~~~~

First system is good to apply the Substitution method.

    X + 3Y = 2      (1)
   2X + 5Y = 3      (2)


From equation (1), express  X = -3Y + 2.  Then substitute it into equation (2),  replacing X there.  
You will get then

    2*(-3Y + 2) + 5Y = 3    (3)


It is equation for one single unknown Y. Simplify it step by step and find Y

    -6Y + 4 + 5Y = 3

    -Y           = 3 - 4 

     -Y          = -1

      Y          = 1.


Now substitute the found value  Y = 1  into equation (1)  to find X.


    X + 3*1 = 2,

or

    X + 3 = 2,

    X     = 2 - 3 = -1.


ANSWER.  The solution is  X= -1,  Y = 1.

Check it  on your own  that the found values  SATISFY  the given equations.

I did this check MENTALLY for myself.


/\/\/\/\/\/\/\/\/

NEVER  POST  more than one problem/question/assignment per post,

otherwise in response you will get a terrible mess,  from which it is  IMPOSSIBLE  to learn.